### v1

# A note on solving basic equations over the semiring of functional digraphs

Abstract
Endowing the set of functional graphs (FGs) with the sum (disjoint union of graphs) and product (standard direct product on graphs) operations induces on FGs a structure of a commutative semiring $R$. The operations on $R$ can be naturally extended to the set of univariate polynomials $R[X]$ over $R$. This paper provides a polynomial time algorithm for deciding if equations of the type $AX=B$ have solutions when $A$ is just a single cycle and $B$ a set of cycles of identical size. We also prove a similar complexity result for some variants of the previous equation.
keywords:
functional graphs, direct product, graphs factorization, equations on graphs
1 Introduction
A functional digraph (FG) is the digraph of a function with finite domain i.e. a digraph with outgoing degree $1$. Endowing the set of functional digraphs FGs with the sum (disjoint union of digraphs) and product (standard direct product on digraphs) operations provides FGs with the structure of a commutative semiring $R$ in which the empty graph (resp., the single loop) is the neutral element of addition (resp., of product) Dennunzio et al. (2018). This semiring can be naturally extended to the semiring of multivariate polynomials $R[X_{1},X_{2},\ldots]$.
A polynomial in $R[X_{1},X_{2},\ldots]$ represents a (infinite) set of functions which have a common substructure provided by the coefficients (modulo isomorphism). The factorization is an interesting inverse problem in this context. Indeed, assume to have a functional graph $G$. Can $G$ be factozised i.e. decomposed into the product or the sum (or a combination of the two) of functional graphs with smaller vertex set [^1] We stress that factorization here is meant in the context of abstract algebra which has nothing to do with the notion of factor graph often used in graph theory.? If this question has been largely investigated for general graphs, little is known for functional digraphs.
Let us consider another version of the factorization problem. Assume that we have partial information (or partial assumptions) on the factorization of $G$. This partial information is represented by a sequence of functional digraphs $A_{1},A_{2},\ldots,A_{k}$ and the question of factorization can be reformulated as `does the following equation
$A_{1}\cdot X_{1}+A_{2}\cdot X_{2}+\ldots A_{k}\cdot X_{k}=G$
admit a solution?'And more generally one can ask for solutions of the following
$A_{1}\cdot X_{1}^{w_{1}}+A_{2}\cdot X_{2}^{w_{2}}+\ldots A_{k}\cdot X_{k}^{w_{k}}=G$ (1)
[…]
$AX=B$ (2)
where $A$ is made by a single cycle of size $p$ and $B$ is a union of cycles of size $q$ (of course, it might be $p\neq q$). In this paper we prove that deciding if equations of type (2) have solutions or not is in ¶.
The interest of this result is manyfold. Without any doubt it will significantly improve the performances of the software pipeline in  Dennunzio et al. (2023). It will shed some light in some connected important questions concerning the cancellation problem for functional graphs i.e. the problem of establishing if $G\cdot F\approx G\cdot H$ implies $F\approx H$ where $F,G$ and $H$ are FGs and $\approx$ is graph isomorphism Doré et al. (2022); Naquin and Gadouleau (2022). We refer the reader to the classical book of Hammak for the factorization and the cancellation problems for more general graphs Hammack et al. (2011).
2 Background and basic facts
In Dennunzio et al. (2018), an abstract algebraic setting for studying finite discrete dynamical systems was introduced. This setting, can be adapted in a straightforward way to FGs. In this section we recall just the minimal concepts to understand the problem we want to solve.
Definition 1 (Sum of FGs).
For any pair of FGs $F=\left(V_{F},E_{F}\right)$ and $G=\left(V_{G},E_{G}\right)$, the sum $H=\left(V_{H},E_{H}\right)$ of $F$ and $G$ is defined as
$V_{H}$ $=V_{F}\sqcup V_{G}$
$E_{H}$ =$=\left\{((a,0),(b,0))\text{ s.t. }(a,b)\in E_{F}\right\}\cup\left\{((a,1),(b,1))\text{ s.t. }(a,b)\in E_{G}\right\}$
where $A\sqcup B$ is the disjoint union of $A$ and $B$ defined as $A\sqcup B=\left(A\times\left\{0\right\}\right)\cup\left(B\times\left\{1\right\}\right)$.
The product of FGs is the standard direct product of graphs that we recall here for completeness sake.
Definition 2 (Product of FGs).
For any pair of FGs $F=\left(V_{F},E_{F}\right)$ and $G=\left(V_{G},E_{G}\right)$, the product $H=\left(V_{H},E_{H}\right)$ of $F$ and $G$ is defined as
$V_{H}$ $=V_{F}\times V_{G}$
$E_{H}$ =$=\left\{\left((a,c),(b,d)\right)\in V_{H}\times V_{H}\text{ s.t }(a,b)\in E_{F}\text{ and }(c,d)\in E_{G}\right\}$
When no misunderstanding is possible we will denote with $+$ (resp., $\cdot$) the sum (resp., the product) of FGs.
As already informally stated in the introduction we have the following algebraic characterisation of the class of FGs.
Theorem 1 (Dennunzio et al. (2018)).
The class of FGs (modulo isomorphisms) equipped with the operations of $+$ and normal-$\cdot$ is a commutative semi-ring.
[…]
Figure 1: An example of polynomial equation over FGs (above) and its solution (below).
From Dennunzio et al. (2018) we also recall the following important result about the complexity of solving polynomial equations.
Theorem 2.
The problem of deciding if there are solutions to
  • 1.
    polynomial equations over $R[X_{1},X_{2},\ldots]$ is undecidable;
  • 2.
[…]
It is therefore natural to ask if simpler equations admit solution algorithms with lower complexity.
3 Solving basic equations on permutations
From now on we will focus on a subclass of functional graphs, namely the graphs of permutations. For this reason, we introduce a convenient notation called C-notation taken from Dennunzio et al. (2023). We note $C^{1}_{p}$ the graph made by a single cycle of size $p$. Similarly, $C^{n}_{p}$ denotes a graph which is the sum of $n$ graphs $C^{1}_{p}$. According to this notation, it is clear that $C^{n_{1}}_{p}+C^{n_{2}}_{p}=C^{n_{1}+n_{2}}_{p}$ and $kC^{n}_{p}=C^{kn}_{p}$ for any natural $n_{1},n_{2}$ and $k$. With this notation, the classical result that the graph $G$ of a permutation is a union of disjoint cycles translates into
$G=\sum_{i=1}^{l}C_{p_{i}}^{n_{i}}$
for suitable positive integers $p_{1},\ldots,p_{l}$ and $n_{1},\ldots,n_{l}$. The following proposition provides an explicit expression for the product of unions of cycles.
Proposition 1.
For any natural $l>1$ and any positive naturals $n_{1},\ldots,n_{l}$, $p_{1},\ldots,p_{l}$, it holds
$\bigodot\limits_{i=1}^{l}C^{n_{i}}_{p_{i}}=C^{\frac{\tilde{p}_{l}\tilde{n}_{l}}{\lambda_{l}}}_{\lambda_{l}}\enspace.$
where $\lambda_{l}=\operatorname{lcm}(p_{1},\ldots,p_{l})$, $\tilde{p}_{l}=\prod_{i=1}^{l}p_{i}$ and $\tilde{n}_{l}=\prod_{i=1}^{l}n_{i}$.
Proof.
We proceed by induction over $l$. First of all, we prove that the statement is true for $l=2$, i.e.,
$C^{n_{1}}_{p_{1}}\cdot C^{n_{2}}_{p_{2}}=C^{\frac{(p_{1}p2)(n_{1}n_{2})}{\lambda_{2}}}_{\lambda_{2}}\enspace.$ (3)
Let us consider the case $n_{1}=n_{2}=1$. Since $C^{1}_{p_{1}}$ and $C^{1}_{p_{2}}$ can be viewed as finite cyclic groups of order $p_{1}$ and $p_{2}$, respectively, each element of the product of such cyclic groups has order $\lambda_{2}=\operatorname{lcm}(p_{1},p_{2})$ or, in other words, each element of $C^{1}_{p_{1}}\odot C^{1}_{p_{2}}$ belongs to some cycle of length $\lambda_{2}$. So, $C^{1}_{p_{1}}\odot C^{1}_{p_{2}}$ just consists of $(p_{1}\cdot p_{2})/\lambda_{2}$ cycles, all of length $\lambda_{2}$, and therefore
$C^{1}_{p_{1}}\cdot C^{1}_{p_{2}}=C^{\frac{p_{1}p_{2}}{\lambda_{2}}}_{\lambda_{2}}\enspace.$
In the case $n_{1}\neq 1$ or $n_{2}\neq 1$, since the product is distributive over the sum, we get
$C^{n_{1}}_{p_{1}}\cdot C^{n_{2}}_{p_{2}}=\sum\limits_{i=1}^{n_{1}}C^{1}_{p_{1}}\cdot\;\sum\limits_{j=1}^{n_{2}}C^{1}_{p_{2}}=\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}(C^{1}_{p_{1}}\cdot\;C^{1}_{p_{2}})=\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}C^{\frac{p_{1}p_{2}}{\lambda_{2}}}_{\lambda_{2}}=C^{\frac{(p_{1}p_{2})(n_{1}n_{2})}{\lambda_{2}}}_{\lambda_{2}}\enspace.$
Assume now that the equality holds for any $l>1$. Then, we get
$\sum\limits_{i=1}^{l+1}C^{n_{i}}_{p_{i}}=C^{\frac{\tilde{p}_{l}\tilde{n}_{l}}{\lambda_{l}}}_{\lambda_{l}}\cdot\;C^{n_{l+1}}_{p_{l+1}}=C^{\frac{\tilde{p}_{l}p_{l+1}\tilde{n}_{l}n_{l+1}}{\operatorname{lcm}(\lambda_{l},p_{l+1})}}_{\operatorname{lcm}(\lambda_{l},p_{l+1})}=C^{\frac{\tilde{p}_{l+1}\tilde{n}_{l+1}}{\lambda_{l+1}}}_{\lambda_{l+1}}\enspace,$
where $\lambda_{l+1}=\lambda(\lambda_{l},p_{l+1})$, $p_{l+1}=\tilde{p}_{l}p_{l+1}$ and $n_{l+1}=\tilde{n}_{l}n_{l+1}$. ∎
According to the $C$-notation basic equations will have the following form
$C^{1}_{p}\cdot X=C^{n}_{q}\enspace,$ (4)
where $X$ is the unknown and $p,n$ and $q$ are positive integers.
As we already stressed in the introduction, in many cases, we are not interested in knowing the exact solutions of a basic equation over permutations but we are interested in deciding if an equation admits a solution or not. More formally, in this paper we are interested in the computational complexity of the following decision problem.
[…]
   Question: Does Equation (4) admit any solution?
[…]
Theorem 3.
Let $p,q,n$ be positive integers. Moreover, assume that the prime factorisations of $p,q$ and $n$ are:
$p$ $=p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}\text{ with }\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$
$q$ $=q_{1}^{k_{1}}\cdot\ldots\cdot q_{r}^{k_{s}}\text{ with }\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$
$n$ $=n_{1}^{m_{1}}\cdot\ldots\cdot n_{s}^{m_{t}}\text{ with }\operatorname{factors}\left(n\right)=\left\{n_{1},\ldots,n_{t}\right\}\enspace.$
The equation $C^{1}_{p}\cdot\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}=C^{n}_{q}$ has solutions if and only if the two following conditions hold
  1. 1.
    $p$ divides $q$;
  2. 2.
    for all $i\in\left\{1,\ldots,r\right\}$ either $p_{i}^{h_{i}}\in\left\{q_{1}^{k_{1}},\ldots,q_{1}^{k_{s}}\right\}$ or $p_{i}^{h_{i}}$ divides $n$.
Proof.
The left part of equation can be rewritten as $\sum_{i=1}^{l}C^{1}_{p}\cdot C^{1}_{p^{\prime}_{i}}$. Hence, by Proposition 1, $p$ must divide $q$. This means that $\operatorname{factors}\left(p\right)\subseteq\operatorname{factors}\left(q\right)$ and that for any $p_{i}^{h_{i}}$ there exists $j\in\left\{1,\ldots,s\right\}$ such that $q_{j}\in\operatorname{factors}\left(q\right)$ and $h_{i}\leq k_{j}$.
($\Leftarrow$) Given the factorisation of $q$ and Proposition 1, we can rewrite $C^{n}_{q}$ as follows.
$C^{n}_{q}=C^{1}_{q}\cdot C^{n}_{1}=C^{1}_{{q_{1}^{k_{1}}}\cdot\ldots\cdot{q_{s}^{k_{s}}}}\,\cdot C^{n}_{1}=C^{1}_{q_{1}^{k_{1}}}\cdot\ldots\cdot C^{1}_{q_{s}^{k_{s}}}\cdot C^{n}_{1}$
Suppose now that for some $i\in\left\{1,\ldots,r\right\}$, ${p_{i}^{h_{i}}}$ is not contained in $\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$. Then, there exists $j\in\left\{1,\ldots,s\right\}$ such that $p_{i}=q_{j}$ and $h_{i}<k_{j}$. Now, if $p_{i}^{h_{i}}$ divides $n$, we have
$C^{1}_{q_{1}^{k_{1}}}\cdot\ldots\cdot C^{1}_{q_{s}^{k_{s}}}\cdot C^{n}_{1}=C^{1}_{q_{1}^{k_{1}}}\cdot\ldots\cdot C^{1}_{q_{s}^{k_{s}}}\cdot C^{1}_{{\mathsf{p}_{i}^{h_{i}}}}\cdot C^{\frac{n}{\left(p_{i}^{h_{i}}\right)}}_{1}.$
By repeating the same operation for all $p_{i}^{h_{i}}$ not included in $\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$, we find a solution since $\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}$ will contain a cycle $C^{1}_{p^{\prime}_{i}}$ (with $p^{\prime}_{i}$ equals $q_{j}^{k_{j}}$) for any $p_{i}^{h_{i}}$ not included, and the set of self-loops.
($\Rightarrow$) Let us suppose that the equation $C^{1}_{p}\cdot\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}=C^{n}_{q}$ holds for some $p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{l}$. Let us assume that there exists $i$ such that $p_{i}^{h_{i}}\notin\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$. Then, $p_{i}^{h_{i}}$ will be equal to a certain $q_{j}^{y}$ with $y<k_{j}$. This implies that all $p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{l}$ must contain exactly $q_{j}^{k_{j}}$ as a factor, since the $\operatorname{lcm}$ with $p$ must be $q$. Thus we can write:
$\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}$ $=C^{1}_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C^{1}_{\frac{p^{\prime}_{i}}{\left(q_{j}^{k_{j}}\right)}}$
which leads us to the following
$C^{1}_{p}\cdot\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}$ $=C^{1}_{p}\cdot C^{1}_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C^{1}_{\frac{p^{\prime}_{i}}{\left(q_{j}^{k_{j}}\right)}}$
$=C^{1}_{\frac{p}{\left(q_{j}^{y}\right)}}\cdot C^{1}_{q_{j}^{y}}\cdot C^{1}_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C^{1}_{\frac{p^{\prime}_{i}}{\left(q_{j}^{k_{j}}\right)}}$
$=C^{1}_{\frac{p}{\left(q_{j}^{y}\right)}}\cdot C^{\left(q_{j}^{y}\right)}_{\left(q_{j}^{k_{j}}\right)}\cdot\sum_{i=1}^{l}C^{1}_{\frac{p^{\prime}_{i}}{\left(q_{j}^{k_{j}}\right)}}$
Then, according to Proposition 1, $p_{i}^{h_{i}}$ (i.e. $q_{j}^{y}$) must divide $n$. ∎
[…]
Lemma 1.
Let $p,q$ be two positive integers. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{s}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{F}$ be the set of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$. Algorithm 1 computes $\Pi_{\mathbf{F}}=\prod_{\mathbf{f}\in\mathbf{F}}\mathbf{f}$ without using the factorizations of $p$ and $q$. Moreover, the time complexity of Algorithm 1 is $O(\log p\cdot\log q)$.
Proof.
Denote $P=\operatorname{factors}\left(p\right)$ and $Q=\operatorname{factors}\left(q\right)$. If $g=\gcd(p,q)=1$ (line $2$), then $P\cap Q=\emptyset$ and hence $\Pi_{\mathbf{F}}=q$. Hence the algorithm returns $q$ (line $3$). If $g\neq 1$, then taking $q/g$ keeps unchanged the $q_{j}^{k_{j}}$ belonging to $Q\setminus P$ and at the same time decreases of at least $1$ the exponents of the $q_{j}$ belonging to $P\cap Q$. Hence, calling recursively (line $5$) $\Pi_{\mathbf{F}}$ with $(q=q/g,g)$ keeps decreasing the exponents of $q_{j}$ belonging to $P\cap Q$ until they become 00. At that point, $\gcd(q,g)=1$ and the algorithm exits returning $\Pi_{\mathbf{F}}$. Remark that $\gcd(p,q)$ can be computed in $O(\log(\min(p,q)))$ and that the depth of the recursion is $O(\log q)$ in the worst case. Hence, the worst case complexity is $O(\log p\cdot\log q)$. ∎
1 Function Compute-$\mathbf{\Pi_{F}}$(p, q)
      Input : $p$ and $q$ positive integers
      Output : $\Pi_{\mathbf{F}}$
2      $g\leftarrow\gcd(p,q)$;
3      if $g$ == $1$ then
4           return q;
5
6           else
7                return ${\Pi}_{\mathbf{F}}(g,\frac{q}{g})$;
8
9                end if
10
11
Algorithm 1
Lemma 2.
Let $p$ and $q$ be two positive integers such that $p|q$. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{s}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{E}$ be the set of all $p_{i}=q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Then, Algorithm 2 computes ${\Pi}_{\mathbf{E}}=\prod_{\mathbf{e}\in\mathbf{E}}\mathbf{e}$ without using the factorizations of $p$ and $q$. The time complexity of Algorithm 2 is $O(\log p\cdot\log\log p)$.
Proof.
By executing lines $2$ and $3$ of Algorithm 2 we save in $g$ the product of the factors $q^{k_{j}}_{j}$ such that $q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$. However, the exponents of the factors $q_{j}$ in $g$ are those of $q$ while to compute $\mathbf{\Pi_{E}}$ we need those of $p$. This is the purpose of the while loop (lines $5$-$9$). Indeed, by recursively squaring $d$ it will make the exponents of the factors of $q$ selected in $g$ grow bigger than those in $p$ and hence the $\gcd$ of line $8$ will select the factors and the corresponding exponents that we were looking for. Concerning the complexity, it is enough to remark that the number of iterations of the while loop is $O(\log\log p)$ in the worst case and that each iteration costs $O(\log(\min(q/p,p)))$. ∎
1 Function Compute-$\mathbf{\Pi_{E}}$(p, q)
      Input : $p$ and $q$ integers s.t. $p|q$
      Output : $\prod_{\mathbf{e}\in\mathbf{E}}\mathbf{e}$
2      $d\leftarrow\frac{q}{p}$;
3      $g\leftarrow\gcd(d,p)$;
4      $g^{\prime}\leftarrow\gcd(d*d,p)$;
5      while $g\neq g^{\prime}$ do
6           $g\leftarrow g^{\prime}$;
7           $d\leftarrow d*d$;
8           $g^{\prime}\leftarrow\gcd(d,p)$;
9
10           end while
11          return $g$;
12
13
Algorithm 2
Theorem 4.
[…]
Proof.
[…]
Concerning the complexity, we have that the $\gcd$ at line $2$ can be computed in $O(\log(\min(p,q)))$. By Lemmas 1 and 2, we know that computing $e$ takes $O(\max(\log p\cdot\log\log p,\log p\cdot\log q))$ that is $O(\log p\cdot\log q)$ since $p$ divides $q$. Finally, the check at Line $6$ has time complexity $O(\log(\min(e,n)))\subseteq O(\log p)$ since $e$ is a divisor of $p$. We conclude that the time complexity is $O(\log p\cdot\log q)$. ∎
1 Function DEEP(p, q, n)
      Input : $p,q$ and $n$ positive integers
      Output : $true$ if $C^{1}_{p}\odot X=C^{n}_{q}$ has solutions, $false$ otherwise
2      if $\gcd(p,q)\neq p$ then   // $p$ does not divide $q$
3           return $false$;
4
5           else     // $p$ divides $q$
6                $e\leftarrow\Pi_{\mathbf{E}}(p,\frac{q}{{\Pi}_{\mathbf{F}}(p,q)})$;
7                if $\gcd(e,n)==e$ then   // $e$ divides $n$
8                     return $true$;
9
10                     else
11                          return $false$;
12
13                          end if
14
15                          end if
16
17
Algorithm 3
Example 1.
Consider the following equation
$C^{1}_{8400}\odot X=C^{6000}_{8316000}$
where $p=8400$, $q=8316000$ and $n=6000$. Let us first see how by means of the presented algorithms that we can verify the conditions of Theorem 3 without knowing the factorisations of the numbers involved. Later, we will see how the method acts at the level of the factorisations.
Since $p$ divides $q$, we want to calculate ${\Pi}_{\mathbf{E}}(\frac{8316000}{{\Pi}_{\mathbf{F}}(8316000,8400)},8400)$. Let us begin by considering ${\Pi}_{\mathbf{F}}(8316000$ ,8400),8400), 8400 ). Since $8316000$ and $8400$ are not coprime, the method is iterated, i.e.
${\Pi}_{\mathbf{F}}(\frac{8316000}{\gcd(8316000,8400)},\gcd(8316000,8400))={\Pi}_{\mathbf{F}}(990,8400).$
Again, $990$ and $8400$ are not coprime, we call recursively the function ${\Pi}_{\mathbf{F}}(\frac{990}{\gcd(990,8400)},\gcd(990,8400))={\Pi}_{\mathbf{F}}(33,30)$ which brings us to ${\Pi}_{\mathbf{F}}(\frac{33}{\gcd(33,30)},\gcd(33,30))={\Pi}_{\mathbf{F}}(11,3)$. Since $11$ and $3$ are coprime, the method returns $11$.
Let us therefore study ${\Pi}_{\mathbf{E}}(\frac{8316000}{11},8400)={\Pi}_{\mathbf{E}}(756000,8400)$. With $i=1$, $div=\frac{756000}{8400}=90$, $g=\gcd(90,8400)$ $=30$ and $g^{\prime}=\gcd(90^{2},8400)=\gcd(8100,8400)=300$. Since $g\neq g^{\prime}$, $i$ becomes $2$, $g$ takes value $300$ and $g^{\prime}$ becomes $\gcd(90^{3},8400)=\gcd(729000,8400)=600$. Hence, since $g$ and $g^{\prime}$ are still not equal, the method continues as follows.
$i=3,g=600,g^{\prime}=\gcd(90^{4},8400)=\gcd(65610000,8400)=1200$
$i=4,g=1200,g^{\prime}=\gcd(90^{5},8400)=\gcd(5904900000,8400)=1200$
At this point, since $g=g^{\prime}$, the method returns $1200$. Finally, since $1200$ divides $n=6000$ we know that the equation admits of solution.
Let us now see what happens, from the point of view of the factorisations, by applying the previous method. Considering the values in input of this example, the factorisations are:
$p=2^{4}\cdot 3\cdot 5^{2}\cdot 7,\;\;q=2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11,\;\;n=2^{4}\cdot 3\cdot 5^{3}$
The goal of ${\Pi}_{\mathbf{F}}$ is to compute the product of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$, and, in fact, the method calculates the following.
${\Pi}_{\mathbf{F}}$ $(2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)=$
$={\Pi}_{\mathbf{F}}(\frac{2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11}{\gcd(2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)},\gcd(2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7))$
$={\Pi}_{\mathbf{F}}(2\cdot 3^{2}\cdot 5\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)$
white
${\Pi}_{\mathbf{F}}$ $(2\cdot 3^{2}\cdot 5\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)=$
$={\Pi}_{\mathbf{F}}(\frac{2\cdot 3^{2}\cdot 5\cdot 11}{\gcd(2\cdot 3^{2}\cdot 5\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)},\gcd(2\cdot 3^{2}\cdot 5\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7))$
$={\Pi}_{\mathbf{F}}(3\cdot 11,2\cdot 3\cdot 5)$
white
${\Pi}_{\mathbf{F}}$ $(3\cdot 11,2\cdot 3\cdot 5)=$
$={\Pi}_{\mathbf{F}}(\frac{3\cdot 11}{\gcd(3\cdot 11,2\cdot 3\cdot 5)},\gcd(3\cdot 11,2\cdot 3\cdot 5))$
$={\Pi}_{\mathbf{F}}(11,3)=11.$
Once ${\Pi}_{\mathbf{F}}$ has been calculated, through $\frac{q}{{\Pi}_{\mathbf{F}}(q,p)}$ we obtain the product of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\cap\operatorname{factors}\left(p\right)$, i.e. $2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7$. Then, the goal of ${\Pi}_{\mathbf{E}}$ is to calculate the product of all $q_{j}^{k_{j}}$ such that $q_{j}=p_{i}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Note that $90=2\cdot 3^{2}\cdot 5$.
$i=1,$ $g=\gcd(2\cdot 3^{2}\cdot 5,2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2\cdot 3\cdot 5,g^{\prime}=\gcd(2^{2}\cdot 3^{4}\cdot 5^{2},2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2^{2}\cdot 3\cdot 5^{2},$
$i=2,$ $g=2^{2}\cdot 3\cdot 5^{2},g^{\prime}=\gcd(2^{3}\cdot 3^{6}\cdot 5^{3},2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2^{3}\cdot 3\cdot 5^{2},$
$i=3,$ $g=2^{3}\cdot 3\cdot 5^{2},g^{\prime}=\gcd(2^{4}\cdot 3^{8}\cdot 5^{4},2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2^{4}\cdot 3\cdot 5^{2},$
$i=4,$ $g=2^{4}\cdot 3\cdot 5^{2},g^{\prime}=\gcd(2^{5}\cdot 3^{10}\cdot 5^{5},2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2^{4}\cdot 3\cdot 5^{2}.$
Corollary 1.
Given four positive integers $p,q,n$ and $m$, the problem of deciding if the following equation
$C_{p}^{m}\cdot X=C_{q}^{n}$ (5)
has solutions has time complexity $O(\max(\log\min(m,n),\log p\cdot\log q))$.
Proof.
Assume that Equation (5) has a solution $\sum_{i=1}^{l}C_{t_{i}}^{s_{i}}$ then we have
$C^{n}_{q}=C_{p}^{m}\cdot\sum_{i=1}^{l}C_{t_{i}}^{s_{i}}=\sum_{i=1}^{l}C_{p}^{m}\cdot C_{t_{i}}^{s_{i}}=\sum_{i=1}^{l}C_{\operatorname{lcm}(p,t_{i})}^{ms_{i}}=C_{q}^{\sum_{i=1}^{l}ms_{i}}=C_{q}^{m\sum_{i=1}^{l}s_{i}}$
and hence Equation (5) has a solution iff $m$ divides $n$ and $C_{p}^{1}\cdot X=C^{n/m}_{q}$ has a solution. We conclude that the time complexity of deciding if Equation (5) has a solution is the max between $O(\log\min(m,n))$ (for checking if $m$ divides $n$) and the time complexity of DEEP with parameters $p,q$ and $n/m$ (see Theorem 4). ∎
Corollary 2.
Given $22s+$ 2 positive integers $p,m,q_{1},\ldots,q_{s},n_{1},\ldots,n_{s}$, the problem of deciding if the following equation
$C_{p}^{m}\cdot X=\sum_{i=1}^{s}C_{q_{i}}^{n_{i}}$ (6)
has solutions has time complexity $O(\max(\log\min(m,n),\log p\cdot\log q))$.
Proof.
It is clear that Equation (6) has solutions iff all the $s$ equations of type $C^{m}_{p}\cdot X=C^{n_{i}}_{q_{i}}$ have solutions. The last condition can be checked by applying a finite number of times Algorithm 3. Hence the time complexity is the same as the one of DEEP. ∎
Proposition 2.
Given $r+1$ positive integers $p_{1},\ldots,p_{r},q$ and $r+1$ integers $m_{1},\ldots,m_{r}$ and $n$. If the following equation
$\sum_{i=1}^{r}C_{p_{i}}^{m_{i}}\cdot X=C_{q}^{n}$ (7)
has solutions then $\sum_{k=1}^{r}m_{k}p_{k}$ must divide $nq$.
Proof.
Denote $d(q)$ the set of divisors of $q$. A generic solution to Equation (7) (if it exists) has the form $\hat{X}=\sum_{i=1}^{|d(q)|}C_{t_{i}}^{s_{i}}$, where $t_{i}$ is the $i$-th divisor of $q$ (divisors are considered ordered by the standard order on integers) and $s_{i}$ is an integer (possibly 00). We have
$C^{n}_{q}=\sum_{k=1}^{r}C_{p_{k}}^{m_{k}}\cdot\sum_{i=1}^{|d(q)|}C_{t_{i}}^{s_{i}}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}C_{p_{k}}^{m_{k}}\cdot C_{t_{i}}^{s_{i}}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}C_{\operatorname{lcm}(p_{k},t_{i})}^{m_{k}s_{i}\gcd(p_{k},t_{i})}$ (8)
Since $\hat{X}$ is a solution it must be either $q=\operatorname{lcm}(p_{k},t_{i})$ or $s_{i}=0$. Hence, from $p_{k}t_{i}=\gcd(p_{k},t_{i})\operatorname{lcm}(p_{k},t_{i})$ and the previous consideration we get $\gcd(p_{k},t_{i})=p_{k}t_{i}/q$. If we replace this last quantity in Equation (8) we get
$C^{n}_{q}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}C_{q}^{m_{k}s_{i}p_{k}t_{i}/q}$ (9)
which holds iff
$\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}m_{k}p_{k}s_{i}t_{i}/q=n$
that is to say iff
$\left(\sum_{k=1}^{r}m_{k}p_{k}\right)\left(\sum_{i=1}^{|d(q)|}s_{i}t_{i}\right)=nq$ (10)
We deduce that if Equation (10) has solutions then $\sum_{k=1}^{r}m_{k}p_{k}$ divides $np$. ∎
Let us make some remarks concerning the proof of the previous results. Assume that $H=\frac{nq}{\sum_{k=1}^{r}m_{k}p_{k}}$ is an integer. Then the Diophantine equation
$\sum_{i=1}^{|d(q)|}s_{i}t_{i}=H$ (11)
has solutions iff $\gcd(t_{1},\ldots,t_{|d(q)|})=q$ divides $H$ which is always true if $\sum_{k=1}^{r}m_{k}p_{k}$ divides $n$. With standard techniques (see for instance (Niven et al., 1991, Theorem 5.1, page 213)), one can recursively find the solutions in logarithmic time in $q$. We point out that it might happen that all the solutions are non admissible i.e. every solution contains at least one $s_{i}<0$ for $i\in\left\{1,\ldots,r\right\}$. This issue can also be settled in logarithmic time in $q$ (see the considerations following Theorem 5.1 in (Niven et al., 1991, pages 213-214)). However, the big problem that we have is that we need to compute the divisors of $q$ and, as far as we know, the best algorithms for doing that have time complexity which is a polynomial in $q$.
In a similar manner as in Proposition 2, one can derive another necessary condition for Equation (7) as provided in the following proposition.
Proposition 3.
Given $r+1$ positive integers $p_{1},\ldots,p_{r},q$ and $r+1$ integers $m_{1},\ldots,m_{r}$ and $n$. If Equation (7) has solutions then $\gcd(m_{1},\ldots,m_{k})$ divides $n$.
Proof.
From Equation (8) in the proof of Proposition 2, it is easy to remark that it holds iff
$\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}m_{k}s_{i}\gcd(p_{k},t_{i})=n\enspace.$ (12)
Now, set $S_{k}=\sum_{i=1}^{|d(q)|}s_{i}\gcd(p_{k},t_{i})$ we have
$\sum_{k=1}^{r}m_{k}S_{k}=n\enspace,$ (13)
where the $S_{k}$ are the unknowns. This last equation has solutions iff $\gcd(m_{1},\ldots,m_{r})$ divides $n$. However, this last condition is only necessary since solutions might be non-admissible. ∎
Again, we would like to stress that solutions to Equation (13) can be found in logarithmic time in $\min{m_{1},\ldots,m_{n}}$ giving raise to $r$ equations
$S_{k}=\sum_{i=1}^{|d(q)|}s_{i}\gcd(p_{k},t_{i})$
which can be also solved in logarithmic time provided that divisors $t_{i}$ of $q$ are known. Hence, once again the problem of finding solutions to Equation (7) can be solved in polynomial time in the value of $q$ (more precisely in $O(\max(q,\gcd(m_{1},\ldots,m_{k})))$). However, for some particular values of $q$, Equation (7) can be solved in an effective way. An example is shown by the following proposition.
Proposition 4.
Given $r+1$ positive integers $p_{1},\ldots,p_{r},q$ and $r+2$ integers $m_{1},\ldots,m_{r},t$ and $n$. Let $R=\left\{1,\ldots,r\right\}$. Assume $q$ is prime and that for $i\in R$ we have $p_{i}\neq q^{t}$. Then, the following equation
$\sum_{i=1}^{r}C_{p_{i}}^{m_{i}}\cdot X=C^{n}_{q^{t}}$ (14)
has solutions iff $\sum_{i=1}^{r}m_{i}p_{i}$ divides $n$ and for $i\in R$, $p_{i}=q^{t_{i}}$ for some integer $t_{i}<t$. Moreover, if a solution exists, then it is unique.
Proof.
()($\Rightarrow$)( ⇒ ) Assume that $\sum_{k=1}^{l}C_{v_{k}}^{u_{k}}$ is a solution for the equation. Then we have
$\sum_{i=1}^{r}C_{p_{i}}^{m_{i}}\cdot\sum_{k=1}^{l}C_{v_{k}}^{u_{k}}=\sum_{i=1}^{r}\sum_{k=1}^{l}C_{p_{i}}^{m_{i}}\cdot C_{v_{k}}^{u_{k}}=\sum_{i=1}^{r}\sum_{k=1}^{l}C_{\operatorname{lcm}(p_{i},v_{k})}^{m_{i}u_{k}\gcd(p_{i},v_{k})}=C^{n}_{q^{t}}$
which implies $\operatorname{lcm}(p_{i},v_{k})=q^{t}$. Since $q$ is prime we deduce that either $p_{i}=q^{t}$ or $v_{k}=q^{t}$. Using the hypothesis we have that $v_{k}=q^{t}$ for all $k\in\left\{1,\ldots,l\right\}$. Hence, the solution can be written as $C_{q^{t}}^{u}$. Now, for each $i\in R$ we have $\gcd(p_{i},q^{t})=p_{i}$. Hence, $C_{q^{t}}^{u}$ is a solution iff the following holds
$\sum_{i=1}^{r}C_{p_{i}}^{m_{i}}\cdot C_{q^{t}}^{u}=\sum_{i=1}^{r}C_{q^{t}}^{m_{i}up_{i}}=C^{\sum_{i=1}^{r}m_{i}up_{i}}_{q^{t}}=C^{n}_{q^{t}}$
which holds iff $\sum_{i=1}^{r}m_{i}up_{i}=u\cdot\sum_{i=1}^{r}m_{i}p_{i}=n$.
()($\Leftarrow$)( ⇐ ) It is not difficult to see that if $\sum_{i=1}^{r}m_{i}p_{i}$ divides $n$, then $C^{\frac{n}{\sum_{i=1}^{r}m_{i}p_{i}}}_{q^{t}}$ is a solution provided that $p_{i}$ divides $q^{t}$ which implies that $p_{i}=q^{t_{i}}$ for some $t_{i}\leq t$.
Concerning the uniqueness of the solution, it is enough to remark that question is equivalent to ask for solutions of a Diophantine equation on a single variable. ∎
4 Conclusion
This paper proves that DEEP  and some of its variants are in ¶. For some other variants we were able to prove only polynomially testable necessary conditions. These results will greatly improve the performances of the software pipeline for solving general polynomial equations over functional graphs proposed in Dennunzio et al. (2023).
Several improvements are possible. The first one is to reproduce similar results for general functional graphs. A second one would investigate effective algorithms to enumerate solutions of basic equations in an effective way i.e. in such a way that producing two successive solutions is polynomial in the size of the first solution.
[…]