### v4

# On solving basic equations over the semiring of functional digraphs

Abstract
Endowing the set of functional digraphs with the sum (disjoint union of digraphs) and product (standard direct product on digraphs) operations induces on FGs a structure of a commutative semiring $R$. The operations on $R$ can be naturally extended to the set of univariate polynomials $R[X]$ over $R$. This paper provides a polynomial time algorithm for deciding if equations of the type $AX=B$ have solutions when $A$ is just a single cycle and $B$ a set of cycles of identical size. We also prove a similar complexity result for some variants of the previous equation.
keywords:
functional digraphs, direct product, graphs factorization, equations on digraphs
1 Introduction
A functional digraph (FG) is the digraph of a function with finite domain i.e., a digraph with outgoing degree $1$. Endowing the set of functional digraphs FGs with the sum (disjoint union of digraphs) and product (standard direct product on digraphs) operations provides FGs with the structure of a commutative semiring $R$ in which the empty graph (resp., the single loop) is the neutral element of addition (resp., of product) [3]. This semiring can be naturally extended to the semiring of multivariate polynomials $R[X_{1},X_{2},\ldots,X_{k}]$.
A polynomial in $R[X_{1},X_{2},\ldots,X_{k}]$ represents a (infinite) set of functions which have a common substructure provided by the coefficients (modulo isomorphism). The factorization is an interesting inverse problem in this context. Indeed, assume to have a functional digraph $G$. Can $G$ be factorized i.e., decomposed into the product or the sum (or a combination of the two) of functional digraphs with smaller vertex set [^1] We stress that factorization here is meant in the context of abstract algebra which has nothing to do with the notion of factor graph often used in graph theory.? If this question has been largely investigated for general graphs, little is known for functional digraphs.
Let us consider another version of the factorization problem. Assume that we have partial information (or partial assumptions) on the factorization of $G$. This partial information is represented by a sequence of functional digraphs $A_{1},A_{2},\ldots,A_{k}$ and the question of factorization can be reformulated as “does the following equation
$A_{1}\cdot X_{1}+A_{2}\cdot X_{2}+\ldots+A_{k}\cdot X_{k}=G$
admit a solution?”  And more generally one can ask for solutions of the following
$A_{1}\cdot X_{1}^{w_{1}}+A_{2}\cdot X_{2}^{w_{2}}+\ldots+A_{k}\cdot X_{k}^{w_{k}}=G$ (1)
[…]
$AX=B$ (2)
where $A$ is made by a single cycle of size $p$ and $B$ is a union of cycles of size $q$ (of course, it might be $p\neq q$). In this paper, we prove that deciding if equations of type (2) have solutions or not is in ¶.
The interest of this result is manyfold. Without any doubt, it will significantly improve the performance of the software pipeline in  [4] for the case in which one is interested to the existence of solutions and not to their enumeration. It will shed some light in some connected important questions concerning the cancellation problem for functional digraphs i.e., the problem of establishing if $G\cdot F\approx G\cdot H$ implies $F\approx H$ where $F,G$ and $H$ are FGs and $\approx$ is the relation of graph isomorphism [5, 8]. We refer the reader to the classical book of Hammak for the factorization and the cancellation problems for more general graphs [7].
2 Background and basic facts
In [3], an abstract algebraic setting for studying finite discrete dynamical systems was introduced. This setting, can be adapted in a straightforward way to FGs. In this section, we recall just the minimal concepts to understand the problem we want to solve.
Definition 1 (Sum of FGs).
For any pair of FGs $F=\left(V_{F},E_{F}\right)$ and $G=\left(V_{G},E_{G}\right)$, the sum $H=\left(V_{H},E_{H}\right)$ of $F$ and $G$ is defined as
$V_{H}$ $=V_{F}\sqcup V_{G}$
$E_{H}$ =$=\left\{((a,0),(b,0))\text{ s.t. }(a,b)\in E_{F}\right\}\cup\left\{((a,1),(b,1))\text{ s.t. }(a,b)\in E_{G}\right\}$
where $A\sqcup B$ is the disjoint union of $A$ and $B$ defined as $A\sqcup B=\left(A\times\left\{0\right\}\right)\cup\left(B\times\left\{1\right\}\right)$.
The product of FGs is the standard direct product of graphs that we recall here for completeness sake.
Definition 2 (Product of FGs).
For any pair of FGs $F=\left(V_{F},E_{F}\right)$ and $G=\left(V_{G},E_{G}\right)$, the product $H=\left(V_{H},E_{H}\right)$ of $F$ and $G$ is defined as
$V_{H}$ $=V_{F}\times V_{G}$
$E_{H}$ =$=\left\{\left((a,c),(b,d)\right)\in V_{H}\times V_{H}\text{ s.t }(a,b)\in E_{F}\text{ and }(c,d)\in E_{G}\right\}$
When no misunderstanding is possible we will denote with $+$ (resp., $\cdot$) the sum (resp., the product) of FGs.
As already informally stated in the introduction, we have the following algebraic characterisation of the class of FGs.
Theorem 1 ([3]).
The class of FGs (modulo isomorphisms) equipped with the operations of $+$ and $\cdot$ is a commutative semiring.
[…]
Figure 1: An example of polynomial equation over FGs (above) and a solution (below).
From [3], we also recall the following important result about the complexity of solving polynomial equations.
Theorem 2.
The problem of deciding if there are solutions to
  • 1.
    polynomial equations over $R[X_{1},X_{2},\ldots]$ is undecidable;
  • 2.
[…]
It is therefore natural to ask if simpler equations admit solution algorithms with lower complexity. This will be very convenient in practical applications. Indeed, in Biology, when analyzing the behavior of a gene regulation network,for example, one can ask if the observed behavior is produced by the network at hand or if it is the result of the cooperation of simpler systems. These questions can be easily translated in terms of equations over FGs. However, gene regulation networks may contain hundreds of genes and hence effective methods are necessary to solve those equations.
3 Solving basic equations on permutations
From now on, we will focus on a subclass of functional digraphs, namely, the digraphs of permutations. For this reason, we introduce a convenient notation called C-notation inspired from [4]. We note $C_{p}$ the graph made by a single cycle of size $p$. Similarly, $n\cdot C_{p}$ (or simply $nC_{p}$) denotes a graph that is the sum of $n$ graphs $C_{p}$. According to this notation, it is clear that $aC_{p}+bC_{p}=(a+b)C_{p}$ and $a\cdot bC_{p}=abC_{p}$ for any natural $a,b$. With this notation, the classical result that the graph $G$ of a permutation is a union of disjoint cycles translates into
$G=\sum_{i=1}^{l}n_{i}C_{p_{i}}$
for suitable positive integers $p_{1},\ldots,p_{l}$ and $n_{1},\ldots,n_{l}$. Figure 2 provides an example of operations on permutation digraphs and their expression through the C-notation.
[…]
Figure 2: Example of operations on FGs of permutations. Using the C-notation, we have $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot(\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C_{3}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}C_{2})\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C_{3}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=C_{6}+2C_{2}$.
The following proposition provides an explicit expression for the product of unions of cycles.
Proposition 1.
For any natural $l>1$ and any positive naturals $n_{1},\ldots,n_{l}$, $p_{1},\ldots,p_{l}$, it holds
$\prod\limits_{i=1}^{l}n_{i}C_{p_{i}}=\frac{\tilde{p}_{l}\tilde{n}_{l}}{\lambda_{l}}C_{\lambda_{l}}\enspace.$
where $\lambda_{l}=\operatorname{lcm}(p_{1},\ldots,p_{l})$, $\tilde{p}_{l}=\prod_{i=1}^{l}p_{i}$ and $\tilde{n}_{l}=\prod_{i=1}^{l}n_{i}$.
Proof.
We proceed by induction over $l$. First of all, we prove that the statement is true for $l=2$, i.e.,
$n_{1}C_{p_{1}}\cdot n_{2}C_{p_{2}}=\frac{(p_{1}p_{2})(n_{1}n_{2})}{\lambda_{2}}C_{\lambda_{2}}\enspace.$ (3)
Consider the case $n_{1}=n_{2}=1$. Since $C_{p_{1}}$ and $C_{p_{2}}$ can be viewed as finite cyclic groups of order $p_{1}$ and $p_{2}$, respectively, each element of the product of such cyclic groups has order $\lambda_{2}=\operatorname{lcm}(p_{1},p_{2})$ or, in other words, each element of $C_{p_{1}}\cdot C_{p_{2}}$ belongs to some cycle of length $\lambda_{2}$. So, $C_{p_{1}}\cdot C_{p_{2}}$ consists only of $(p_{1}\cdot p_{2})/\lambda_{2}$ cycles, all of length $\lambda_{2}$, and therefore
$C_{p_{1}}\cdot C_{p_{2}}=\frac{p_{1}p_{2}}{\lambda_{2}}C_{\lambda_{2}}\enspace.$
In the case $n_{1}\neq 1$ or $n_{2}\neq 1$, since the product is distributive over the sum, we get
$n_{1}C_{p_{1}}\cdot n_{2}C_{p_{2}}=\sum\limits_{i=1}^{n_{1}}C_{p_{1}}\cdot\;\sum\limits_{j=1}^{n_{2}}C_{p_{2}}=\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}(C_{p_{1}}\cdot\;C_{p_{2}})=\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}\frac{p_{1}p_{2}}{\lambda_{2}}C_{\lambda_{2}}=\frac{(p_{1}p_{2})(n_{1}n_{2})}{\lambda_{2}}C_{\lambda_{2}}\enspace.$
Assume now that the equality holds for any $l>2$. Then, we get
$\sum\limits_{i=1}^{l+1}C^{n_{i}}_{p_{i}}=\frac{\tilde{p}_{l}\tilde{n}_{l}}{\lambda_{l}}C_{\lambda_{l}}\cdot\;n_{l+1}C_{p_{l+1}}=\frac{\tilde{p}_{l}p_{l+1}\tilde{n}_{l}n_{l+1}}{\operatorname{lcm}(\lambda_{l},p_{l+1})}C_{\operatorname{lcm}(\lambda_{l},p_{l+1})}=\frac{\tilde{p}_{l+1}\tilde{n}_{l+1}}{\lambda_{l+1}}C_{\lambda_{l+1}}\enspace,$
where $\lambda_{l+1}=\operatorname{lcm}(\lambda_{l},p_{l+1})$, $\tilde{p}_{l+1}=\tilde{p}_{l}p_{l+1}$ and $\tilde{n}_{l+1}=\tilde{n}_{l}n_{l+1}$. ∎
In the sequel, we will also make use of the following notation. Let $n$ be any integer strictly greater than $1$. We note $\operatorname{factors}\left(n\right)$ the set of prime factors of $n$ i.e., if $n=n_{1}^{m_{1}}\cdot\ldots\cdot n_{t}^{m_{t}}$, then $\operatorname{factors}\left(n\right)=\left\{n_{1},\ldots,n_{t}\right\}$ where $n_{1},\ldots,n_{t}$ are distinct primes.
According to the $C$-notation, basic equations will have the following form
$C_{p}\cdot X=nC_{q}\enspace,$ (4)
where $X$ is the unknown and $p,n$ and $q$ are positive integers. We stress that the numbers $p,q$ and $n$ are just positive integers. They should not be thought of as prime integers unless explicitly mentioned. One of the issues of basic equations is that they might admit several distinct solutions. This is because some digraphs admit several distinct factorizations. Figure 3 provides an easy example of this fact.
[…]
Figure 3: Example of functional digraph having multiple factorizations. Using the C-notation, we have $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}C_{1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}C_{1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})\cdot\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=2C_{2}$.
As we already stressed in the introduction, in many cases, we are not interested in knowing the exact solutions of a basic equation over permutations but we are interested in deciding if an equation admits a solution or not. More formally, in this paper, we are interested in the computational complexity of the following decision problem.
[…]
   Question: Does the equation $C_{p}\cdot X=nC_{q}$ admit any solution?
[…]
However, we are going to show that the problem can be solved in polynomial time in Theorem 4. The proof of Theorem 4 is based on the following characterization of the solutions of basic equations on the digraphs of permutations.
Theorem 3.
Let $p,q,n$ be positive integers and let the prime decompositions of $p$ and $q$ be as follows:
$p$ $=p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}\text{ with }\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$
$q$ $=q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}\text{ with }\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$
The equation $C_{p}\cdot\sum_{i=1}^{l}C_{z_{i}}=nC_{q}$ has solutions (the values $z_{i}$ are the unknowns and need not to be distinct) if and only if the two following conditions hold:
  1. 1.
    $p$ divides $q$;
  2. 2.
    $e$ divides $n$,
where $e=\prod_{v\in E}v$ with $E=\left\{q_{j}^{k_{j}}\;|\;\exists i\text{\ s.t.\@\ }p_{i}=q_{j}\text{ and }h_{i}<k_{j}\right\}$.
Proof.
The left part of equation can be rewritten as $\sum_{i=1}^{l}C_{p}\cdot C_{z_{i}}$. Hence, by Proposition 1, $p$ must divide $q$. This means that $\operatorname{factors}\left(p\right)\subseteq\operatorname{factors}\left(q\right)$ and that for any $p_{i}^{h_{i}}$ there exists $j\in\left\{1,\ldots,s\right\}$ such that $q_{j}\in\operatorname{factors}\left(q\right)$ and $h_{i}\leq k_{j}$.
($\Leftarrow$) Given the factorization of $q$ and Proposition 1, we can rewrite $nC_{q}$ as follows.
$nC_{q}=C_{q}\cdot nC_{1}=C_{{q_{1}^{k_{1}}}\cdot\ldots\cdot{q_{s}^{k_{s}}}}\,\cdot nC_{1}=C_{q_{1}^{k_{1}}}\cdot\ldots\cdot C_{q_{s}^{k_{s}}}\cdot nC_{1}$
Suppose now that, for some $i\in\left\{1,\ldots,r\right\}$, ${p_{i}^{h_{i}}}$ is not contained in $\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$. Then, there exists $j\in\left\{1,\ldots,s\right\}$ such that $p_{i}=q_{j}$ and $h_{i}<k_{j}$. Now, if $p_{i}^{h_{i}}$ divides $n$, we have
$C_{q_{1}^{k_{1}}}\cdot\ldots\cdot C_{q_{s}^{k_{s}}}\cdot nC_{1}=C_{q_{1}^{k_{1}}}\cdot\ldots\cdot C_{q_{s}^{k_{s}}}\cdot C_{{\mathsf{p}_{i}^{h_{i}}}}\cdot\frac{n}{\left(p_{i}^{h_{i}}\right)}C_{1}.$
Repeating the same operation for all $p_{i}^{h_{i}}$ not included in $\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$, we find a solution since $\sum_{i=1}^{l}C_{z_{i}}$ will contain a cycle $C_{z_{i}}$ (with $z_{i}$ equal to $q_{j}^{k_{j}}$) for any $p_{i}^{h_{i}}$ not included, and the set of self-loops.
($\Rightarrow$) Let us suppose that the equation $C_{p}\cdot\sum_{i=1}^{l}C_{z_{i}}=nC_{q}$ holds for some $z_{1},z_{2},\ldots,z_{l}$. Let us assume that there exists $i$ such that $p_{i}^{h_{i}}\notin\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$. Then, $p_{i}^{h_{i}}$ will be equal to a certain $q_{j}^{y}$ with $y<k_{j}$. This implies that all $z_{1},z_{2},\ldots,z_{l}$ must contain exactly $q_{j}^{k_{j}}$ as a factor, since the $\operatorname{lcm}$ with $p$ must be $q$. Thus we can write:
$\sum_{i=1}^{l}C_{z_{i}}$ $=C_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C_{\frac{z_{i}}{\left(q_{j}^{k_{j}}\right)}}$
which leads us to the following
$C_{p}\cdot\sum_{i=1}^{l}C_{z_{i}}$ $=C_{p}\cdot C_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C_{\frac{z_{i}}{\left(q_{j}^{k_{j}}\right)}}$
$=C_{\frac{p}{\left(q_{j}^{y}\right)}}\cdot C_{q_{j}^{y}}\cdot C_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C_{\frac{z_{i}}{\left(q_{j}^{k_{j}}\right)}}$
$=C_{\frac{p}{\left(q_{j}^{y}\right)}}\cdot q_{j}^{y}C_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C_{\frac{z_{i}}{\left(q_{j}^{k_{j}}\right)}}$
Then, according to Proposition 1, $p_{i}^{h_{i}}$ (i.e., $q_{j}^{y}$) must divide $n$. ∎
Theorem 3 provides a condition to decide whether a solution to a basic equation can exist. However, we are not going to exploit it directly, since it requires the prime factorization of the integers $p$ and $q$ which would imply unnecessarily high complexity. In fact, the algorithms we are going to conceive will essentially use simple arithmetical operations or basic functions on the integers. In this regard, we consider that the worst-case time complexity of the addition (resp., multiplication) between two integers is $O(n)$ (resp., $O(n^{2})$), where $n$ is the input size in bits. The division between two integers is assumed to have the same complexity as the multiplication. Concerning $\gcd$, we consider the classical Euclidean GCD algorithm which has worst-case time complexity $O(n^{2})$.
[…]
Lemma 1.
Let $p,q$ be two positive integers. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{F}$ be the set of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$. Algorithm 1 computes $\Pi_{\mathbf{F}}=\prod_{\mathbf{f}\in\mathbf{F}}\mathbf{f}$ without using the factorizations of $p$ and $q$. Moreover, the worst-case time complexity of Algorithm 1 is $O(n^{3})$, where $n$ is the size of the input in bits.
Proof.
Denote $P=\operatorname{factors}\left(p\right)$ and $Q=\operatorname{factors}\left(q\right)$. If $g=\gcd(p,q)=1$ (Line $2$), then $P\cap Q=\emptyset$ and $\Pi_{\mathbf{F}}=q$. Hence, the algorithm returns $q$ (Line $4$). If $g\neq 1$, then taking $q/g$ keeps unchanged the $q_{j}^{k_{j}}$ belonging to $Q\setminus P$ and at the same time decreases of at least $1$ the exponents of the $q_{j}$ belonging to $P\cap Q$. Therefore, calling recursively (Line $6$) $\Pi_{\mathbf{F}}$ with $(q=q/g,g)$ keeps decreasing the exponents of $q_{j}$ belonging to $P\cap Q$ until they become 00. At that point, $\gcd(q,g)=1$ and the algorithm exits returning $\Pi_{\mathbf{F}}$. Remark that at each call of $\Pi_{F}$, we compute a $\gcd$ (Line $2$) plus a division (Line $6$) which cost $O(n^{2})$ in total. Since the depth of the recursion is $O(n)$, the worst-case time complexity for the algorithm is $O(n^{3})$. ∎
1 Function Compute-$\mathbf{\Pi_{F}}$(p, q)
      Input : $p$ and $q$ positive integers
      Output : $\Pi_{\mathbf{F}}$
2      $g\leftarrow\gcd(p,q)$;
3      if $g$ == $1$ then
4           return q;
5
6           else
7                return ${\Pi}_{\mathbf{F}}(g,\frac{q}{g})$;
8
9                end if
10
11
Algorithm 1
Lemma 2.
Let $p$ and $q$ be two positive integers such that $p|q$. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{E}$ be the set of all $p_{i}=q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Then, Algorithm 2 computes ${\Pi}_{\mathbf{E}}=\prod_{\mathbf{e}\in\mathbf{E}}\mathbf{e}$ without using the factorizations of $p$ and $q$. The worst-case time complexity of Algorithm 2 is $O(n^{2}\log n)$.
Proof.
By executing Lines $2$ and $3$ of Algorithm 2, we save in $g$ the product of the factors $q^{k_{j}}_{j}$ such that $q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$. However, the exponents of the factors $q_{j}$ in $g$ are those of $q$ while to compute $\mathbf{\Pi_{E}}$ we need those of $p$. This is the purpose of the while loop (Lines $5$-$9$). Indeed, by recursively squaring $d$, it will make the exponents of the factors of $q$ selected in $g$ grow bigger than those in $p$ and hence the $\gcd$ of line $8$ will select the factors and the corresponding exponents that we were looking for. Concerning the complexity, it is enough to remark that initialisation part of the algorithm (Lines $2$ to $4$) costs $O(n^{2})$ and that the number of iterations of the while loop is $O(\log n)$ in the worst-case with a cost of $O(n^{2})$ per each iteration. ∎
1 Function Compute-$\mathbf{\Pi_{E}}$(p, q)
      Input : $p$ and $q$ integers s.t. $p|q$
      Output : $\prod_{\mathbf{e}\in\mathbf{E}}\mathbf{e}$
2      $d\leftarrow\frac{q}{p}$;
3      $g\leftarrow\gcd(d,p)$;
4      $g^{\prime}\leftarrow\gcd(d*d,p)$;
5      while $g\neq g^{\prime}$ do
6           $g\leftarrow g^{\prime}$;
7           $d\leftarrow d*d$;
8           $g^{\prime}\leftarrow\gcd(d,p)$;
9
10           end while
11          return $g$;
12
13
Algorithm 2
Theorem 4.
[…]
Proof.
[…]
Concerning the complexity, we have that the $\gcd$ in Line $2$ can be computed in $O(s^{2})$. By Lemmata 1 and 2, we know that computing $e$ takes $O(\max(s^{3},s^{2}\log s,s^{2}))$ that is $O(s^{3})$. Remark that the size in bits of the quantities involved in the calculation of $e$ is bounded by $s$. Indeed, both $\Pi_{F}$ and $\Pi_{E}$ return a divisor of $q$.
Finally, the checks in Lines $2$ and $6$ have complexity $O(s^{2})$ since $e$ is a divisor of $p$ and, in its turn, $p$ is a divisor of $q$. We conclude that the time complexity is $O(s^{3})$. ∎
1 Function DEEP(p, q, n)
      Input : $p,q$ and $n$ positive integers
      Output : $true$ if $C_{p}\cdot X=nC_{q}$ has solutions, $false$ otherwise
2      if $p$ does not divide $q$ then
3           return $false$;
4
5           else     // $p$ divides $q$
6                $e\leftarrow\Pi_{\mathbf{E}}(p,\frac{q}{{\Pi}_{\mathbf{F}}(p,q)})$;
7                if $e$ divides $n$ then
8                     return $true$;
9
10                     else
11                          return $false$;
12
13                          end if
14
15                          end if
16
17
Algorithm 3
Example 1.
Consider the following equation
$C_{8400}\cdot X=6000C_{8316000}$
where $p=8400$, $q=8316000$ and $n=6000$. Let us first see how by means of the presented algorithms that we can verify the conditions of Theorem 3 without knowing the factorizations of the numbers involved. Later, we will see how the method acts at the level of the factorizations.
Since $p$ divides $q$, we want to calculate ${\Pi}_{\mathbf{E}}(\frac{8316000}{{\Pi}_{\mathbf{F}}(8316000,8400)},8400)$. Let us begin by considering ${\Pi}_{\mathbf{F}}(8316000$ ,8400),8400), 8400 ). Since $8316000$ and $8400$ are not coprime, the method is iterated, i.e.,
${\Pi}_{\mathbf{F}}(\frac{8316000}{\gcd(8316000,8400)},\gcd(8316000,8400))={\Pi}_{\mathbf{F}}(990,8400).$
Again, $990$ and $8400$ are not coprime, we call recursively the function ${\Pi}_{\mathbf{F}}(\frac{990}{\gcd(990,8400)},\gcd(990,8400))={\Pi}_{\mathbf{F}}(33,30)$ which brings us to ${\Pi}_{\mathbf{F}}(\frac{33}{\gcd(33,30)},\gcd(33,30))={\Pi}_{\mathbf{F}}(11,3)$. Since $11$ and $3$ are coprime, the method returns $11$.
Let us therefore study ${\Pi}_{\mathbf{E}}(\frac{8316000}{11},8400)={\Pi}_{\mathbf{E}}(756000,8400)$. With $i=1$, $d=\frac{756000}{8400}=90$, $g=\gcd(90,8400)$ $=30$ and $g^{\prime}=\gcd(90^{2},8400)=\gcd(8100,8400)=300$. Since $g\neq g^{\prime}$, $i$ becomes $2$, $g$ takes value $300$ and $g^{\prime}$ becomes $\gcd(90^{3},8400)=\gcd(729000,8400)=600$. Hence, since $g$ and $g^{\prime}$ are still not equal, the method continues as follows.
$i=3,g=600,g^{\prime}=\gcd(90^{4},8400)=\gcd(65610000,8400)=1200$
$i=4,g=1200,g^{\prime}=\gcd(90^{5},8400)=\gcd(5904900000,8400)=1200$
At this point, since $g=g^{\prime}$, the method returns $1200$. Finally, since $1200$ divides $n=6000$ we know that the equation admits a solution.
Let us now see what happens, from the point of view of the factorizations, by applying the previous method. Considering the values in input of this example, the factorizations are:
$p=2^{4}\cdot 3\cdot 5^{2}\cdot 7,\;\;q=2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11,\;\;n=2^{4}\cdot 3\cdot 5^{3}$
so we have $\operatorname{factors}\left(p\right)=\left\{2,3,5,7\right\}$ and $\operatorname{factors}\left(q\right)=\left\{2,3,5,7,11\right\}$. The goal of ${\Pi}_{\mathbf{F}}$ is to compute the product of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$, and, in fact, the method calculates the following.
${\Pi}_{\mathbf{F}}$ $(2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)=$
$={\Pi}_{\mathbf{F}}(\frac{2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11}{\gcd(2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)},\gcd(2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7))$
$={\Pi}_{\mathbf{F}}(2\cdot 3^{2}\cdot 5\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)$
white
${\Pi}_{\mathbf{F}}$ $(2\cdot 3^{2}\cdot 5\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)=$
$={\Pi}_{\mathbf{F}}(\frac{2\cdot 3^{2}\cdot 5\cdot 11}{\gcd(2\cdot 3^{2}\cdot 5\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7)},\gcd(2\cdot 3^{2}\cdot 5\cdot 11,2^{4}\cdot 3\cdot 5^{2}\cdot 7))$
$={\Pi}_{\mathbf{F}}(3\cdot 11,2\cdot 3\cdot 5)$
white
${\Pi}_{\mathbf{F}}$ $(3\cdot 11,2\cdot 3\cdot 5)=$
$={\Pi}_{\mathbf{F}}(\frac{3\cdot 11}{\gcd(3\cdot 11,2\cdot 3\cdot 5)},\gcd(3\cdot 11,2\cdot 3\cdot 5))$
$={\Pi}_{\mathbf{F}}(11,3)=11.$
Once ${\Pi}_{\mathbf{F}}$ has been calculated, through $\frac{q}{{\Pi}_{\mathbf{F}}(q,p)}$ we obtain the product of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\cap\operatorname{factors}\left(p\right)$, i.e., $2^{5}\cdot 3^{3}\cdot 5^{3}\cdot 7$. Then, the goal of ${\Pi}_{\mathbf{E}}$ is to calculate the product of all $q_{j}^{k_{j}}$ such that $q_{j}=p_{i}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Note that $90=2\cdot 3^{2}\cdot 5$.
$i=1,$ $g=\gcd(2\cdot 3^{2}\cdot 5,2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2\cdot 3\cdot 5,g^{\prime}=\gcd(2^{2}\cdot 3^{4}\cdot 5^{2},2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2^{2}\cdot 3\cdot 5^{2},$
$i=2,$ $g=2^{2}\cdot 3\cdot 5^{2},g^{\prime}=\gcd(2^{3}\cdot 3^{6}\cdot 5^{3},2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2^{3}\cdot 3\cdot 5^{2},$
$i=3,$ $g=2^{3}\cdot 3\cdot 5^{2},g^{\prime}=\gcd(2^{4}\cdot 3^{8}\cdot 5^{4},2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2^{4}\cdot 3\cdot 5^{2},$
$i=4,$ $g=2^{4}\cdot 3\cdot 5^{2},g^{\prime}=\gcd(2^{5}\cdot 3^{10}\cdot 5^{5},2^{4}\cdot 3\cdot 5^{2}\cdot 7)=2^{4}\cdot 3\cdot 5^{2}.$
The proof of the following proposition is trivial since it boils down to remark the the number of the vertices of the digraphs involved in the left part of the equation must equal the number of vertices of digraphs in the right part. However, we shall provide an alternative proof to explicit some relations that are discussed right after the end of the proof.
Proposition 2.
Given $r+1$ positive integers $p_{1},\ldots,p_{r},q$ and $r+1$ integers $m_{1},\ldots,m_{r}$ and $n$. If the following equation
$\sum_{i=1}^{r}m_{i}C_{p_{i}}\cdot X=nC_{q}$ (5)
has solutions then $\sum_{i=1}^{r}m_{i}p_{i}$ must divide $nq$.
Proof.
Denote $d(q)$ the set of divisors of $q$. A generic solution to Equation (5) (if it exists) has the form $\hat{X}=\sum_{i=1}^{|d(q)|}s_{i}C_{t_{i}}$, where $t_{i}$ is the $i$-th divisor of $q$ (divisors are considered ordered by the standard order on integers) and $s_{i}$ is an integer (possibly 00). We have
$nC_{q}=\sum_{k=1}^{r}m_{k}C_{p_{k}}\cdot\sum_{i=1}^{|d(q)|}s_{i}C_{t_{i}}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}m_{k}C_{p_{k}}\cdot s_{i}C_{t_{i}}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}m_{k}s_{i}\gcd(p_{k},t_{i})C_{\operatorname{lcm}(p_{k},t_{i})}$ (6)
Since $\hat{X}$ is a solution it must be either $q=\operatorname{lcm}(p_{k},t_{i})$ or $s_{i}=0$. Hence, from $p_{k}t_{i}=\gcd(p_{k},t_{i})\operatorname{lcm}(p_{k},t_{i})$ and the previous consideration, we get $\gcd(p_{k},t_{i})=p_{k}t_{i}/q$. If we replace this last quantity in Equation (6) we get
$nC_{q}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}\frac{m_{k}s_{i}p_{k}t_{i}}{q}C_{q}$ (7)
which holds iff
$\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}m_{k}p_{k}s_{i}t_{i}/q=n$
that is to say iff
$\left(\sum_{k=1}^{r}m_{k}p_{k}\right)\left(\sum_{i=1}^{|d(q)|}s_{i}t_{i}\right)=nq$ (8)
We deduce that if Equation (8) has solutions then $\sum_{k=1}^{r}m_{k}p_{k}$ divides $nq$. ∎
Let us make some remarks concerning the proof of the previous result. Assume that $H=\frac{nq}{\sum_{k=1}^{r}m_{k}p_{k}}$ is an integer. Then, the Diophantine equation
$\sum_{i=1}^{|d(q)|}s_{i}t_{i}=H$ (9)
has solutions iff $\gcd(t_{1},\ldots,t_{|d(q)|})$ divides $H$, which is always true if $\sum_{k=1}^{r}m_{k}p_{k}$ divides $n$. With standard techniques (see for instance [9, Theorem 5.1, page 213]), one can recursively find the solutions in linear time in the size of $q$ in bits. We point out that it might happen that all the solutions are non admissible i.e., every solution contains at least one $s_{i}<0$ for $i\in\left\{1,\ldots,r\right\}$. This issue can also be settled in linear time in the size of $q$ in bits (see the considerations following Theorem 5.1 in [9, pages 213-214]). However, the big problem that we have is that we need to compute the divisors of $q$ and, as far as we know, the best algorithms for computing them have worst-case time complexity which is sub-exponential in the size of $q$ (see [1] for details).
In a similar manner as in Proposition 2, one can derive another necessary condition for Equation (5) as provided in the following proposition.
Proposition 3.
Given $r+1$ positive integers $p_{1},\ldots,p_{r},q$ and $r+1$ integers $m_{1},\ldots,m_{r}$ and $n$. If Equation (5) has solutions then $\gcd(m_{1},\ldots,m_{k})$ divides $n$.
Proof.
Denote $d(q)$ the set of divisors of $q$. From Equation (6) in the proof of Proposition 2, it is easy to check that $\hat{X}=\sum_{i=1}^{|d(q)|}s_{i}C_{t_{i}}$ is a solution if and only if
$\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}m_{k}s_{i}\gcd(p_{k},t_{i})=n\enspace.$ (10)
Now, set $S_{k}=\sum_{i=1}^{|d(q)|}s_{i}\gcd(p_{k},t_{i})$ we have
$\sum_{k=1}^{r}m_{k}S_{k}=n\enspace,$ (11)
where the $S_{k}$ are the unknowns. This last equation has solutions iff $\gcd(m_{1},\ldots,m_{r})$ divides $n$. However, this last condition is only necessary since solutions might be non-admissible. ∎
Again, we would like to stress that solutions to Equation (11) can be found in quadratic time in in the size of the input giving raise to $r$ equations
$S_{k}=\sum_{i=1}^{|d(q)|}s_{i}\gcd(p_{k},t_{i})$
which can be also solved in quadratic time provided that divisors $t_{i}$ of $q$ are known. Hence, once again the problem of finding solutions to Equation (5) can be solved in polynomial time in the size of $q$ in bits (of course, always knowning the divisors of $q$). However, for some particular values of $q$, Equation (5) can be solved in a more effective way. An example is shown by the following proposition.
Proposition 4.
Given $r+1$ positive integers $p_{1},\ldots,p_{r},q$ and $r+2$ integers $m_{1},\ldots,m_{r},t$ and $n$. Let $R=\left\{1,\ldots,r\right\}$. Assume $q$ is prime and that for every $i\in R$ we have $p_{i}\neq q^{t}$. Then, the following equation
$\sum_{i=1}^{r}m_{i}C_{p_{i}}\cdot X=nC_{q^{t}}$ (12)
has solutions iff $\sum_{i=1}^{r}m_{i}p_{i}$ divides $n$ and for every $i\in R$, $p_{i}=q^{t_{i}}$ for some integer $t_{i}<t$. Moreover, if a solution exists, then it is unique.
Proof.
()($\Rightarrow$)( ⇒ ) Assume that $\sum_{k=1}^{l}C_{v_{k}}^{u_{k}}$ is a solution for the equation. Then, we have
$\sum_{i=1}^{r}m_{i}C_{p_{i}}\cdot\sum_{k=1}^{l}u_{k}C_{v_{k}}=\sum_{i=1}^{r}\sum_{k=1}^{l}m_{i}C_{p_{i}}\cdot u_{k}C_{v_{k}}=\sum_{i=1}^{r}\sum_{k=1}^{l}m_{i}u_{k}\gcd(p_{i},v_{k})C_{\operatorname{lcm}(p_{i},v_{k})}=nC_{q^{t}}$
which implies $\operatorname{lcm}(p_{i},v_{k})=q^{t}$. Since $q$ is prime, we deduce that either $p_{i}=q^{t}$ or $v_{k}=q^{t}$. Using the hypothesis we have that $v_{k}=q^{t}$ for all $k\in\left\{1,\ldots,l\right\}$. Hence, the solution can be written as $uC_{q^{t}}$. Now, for every $i\in R$, we have $\gcd(p_{i},q^{t})=p_{i}$. Hence, $uC_{q^{t}}$ is a solution iff the following holds
$\sum_{i=1}^{r}m_{i}C_{p_{i}}\cdot uC_{q^{t}}=\sum_{i=1}^{r}m_{i}up_{i}C_{q^{t}}=\left(\sum_{i=1}^{r}m_{i}up_{i}\right)C_{q^{t}}=nC_{q^{t}}$
which holds iff $\sum_{i=1}^{r}m_{i}up_{i}=u\cdot\sum_{i=1}^{r}m_{i}p_{i}=n$.
()($\Leftarrow$)( ⇐ ) It is not difficult to see that if $\sum_{i=1}^{r}m_{i}p_{i}$ divides $n$, then $\frac{n}{\sum_{i=1}^{r}m_{i}p_{i}}C_{q^{t}}$ is a solution provided that $p_{i}$ divides $q^{t}$ which implies that $p_{i}=q^{t_{i}}$ for some $t_{i}\leq t$.
Concerning the uniqueness of the solution, it is enough to remark that the question is equivalent to asking for solutions of a Diophantine equation on a single variable. ∎
Clearly, the complexity of the algorithm verifying the conditions of Propositions 2, 3 and 4 is quadratic in the size of the input.
4 Conclusions
This paper is concerned with DEEP i.e., the problem of deciding if a basic equation on permutations has a solution or not. We show that DEEP has cubic complexity. However, it is clear that with more care in the choice of the implementation of some of the components of the decision algorithm, we could obtain a time complexity located between $O(n^{1+\epsilon}\log^{k}n)$ and $O(n^{3})$. Our purpose was essentially to prove that the problem could be solved in polynomial time with a polynomial of reasonable degree, but it is clear that when the instances have large sizes, like in some practical applications, then the shift towards more complex implementations of the components should be taken into account.
For some other variants of DEEP, we were able to prove only some necessary conditions which are testable in quadratic time. However, they have some interest of their own as they might also be of help in pruning the search space for software that aims to solve general polynomial equations on functional digraphs such as the software pipeline proposed in [4].
Several extensions to our results are possible. The most natural one consists of moving from functional digraphs of permutations to larger classes of functional digraphs taking into account the new recent results of [5, 6] and [8].
[…]
Acknowledgements
This work was partially supported by:
  • 1.
    the HORIZON-MSCA-2022-SE-01 project 101131549 “Application-driven Challenges for Automata Networks and Complex Systems (ACANCOS)”;
  • 2.
    the PRIN 2022 PNRR project “Cellular Automata Synthesis for Cryptography Applications (CASCA)” (P2022 MPFRT) funded by the European Union – Next Generation EU;
  • 3.
    ANR “Ordinal Time Computations” (OTC), ANR-24-CE48-0335-01.